OpenStudy (abhisar):

What is the component of (3i + 4j) along (i+j) ?

OpenStudy (abhisar):

@hartnn

OpenStudy (abhisar):

@ganeshie8

OpenStudy (abhisar):

@jim_thompson5910

hartnn (hartnn):

component of vector A along B is just \(\Large \dfrac{A.B}{|B|}\)

hartnn (hartnn):

its actually \( |A| \cos \theta \) but we know form the definition of dot product that \(A.B = |A||B| \cos \theta \)

OpenStudy (abhisar):

ok here it is already in vector form, so i did it like this, (3i + 4j).(i+j)= 3+4=7

hartnn (hartnn):

right, thats your numerator

OpenStudy (abhisar):

and now magnitude of (i+j) = root 2

OpenStudy (abhisar):

so answer should be

OpenStudy (abhisar):

??

hartnn (hartnn):

that is correct :)

OpenStudy (abhisar):

but it's not the answer which i should get ! \(\color{green}{\huge\ddot\frown}\)

hartnn (hartnn):

whats the answer that u should get ?

OpenStudy (abhisar):

I think component means vector projection. The formula we used is for scalar projection.

OpenStudy (abhisar):